# AppReferences

### v2

@@ 5 Assessing the Impact of Environmental Policies on Particulate Pollution and Health
As recently as 7 February 2024 the US Environmental Protection Agency (EPA) has proposed a revision of the National Ambient Air Quality Standards (NAAQS) (U.S. Environmental Protection Agency, 2024), which will result in a more stringent environmental policy for the concentration of air pollutants in the US that cause health hazards. To gather information and better understand the implementation of this revised policy, it is critically important to evaluate how past air quality regulations benefitted health outcomes and how these effects varied between different communities.
In our application, we explore this relevant research question by investigating the repercussions of a previous revision of NAAQS, leveraging insights gleaned from historical data and health data. First, we evaluate the direct and indirect effects of the 2005 NAAQS revision on the mortality rate in the period 2010-2016, considering the variation of levels pm2.5, under the principal stratification framework. Second, we take advantage of the flexible Bayesian non-parametric mixture model for the post-treatment variable (that is, pm2.5) to understand how these effects can vary within strata. Third, we characterize the different identified groups of individuals.
@@ 5.1 Data and Study Design
To answer our research question, we have merged two datasets: (i) one dataset containing the information about the NAAQS designations, as well as air pollution levels and the demographic and socioeconomic characteristics in the counties in the eastern U.S., used in the Zigler et al. (2018)’s analysis (data can be found at Zigler, 2017), and (ii) one dataset containing the information on the age-adjusted mortality rate in these counties available by the Center for Disease Control and Prevention of U.S. (Friede et al., 1993).
@@ 5.2 Results
We apply our proposed model to the 384 counties in the eastern US described above, including all covariates, census and meteorological variables, in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 115 counties ($30\%$ of the total counties analyzed), the associative positive stratum with 48 counties ($12.5\%$), and the associative negative stratum with 221 counties ($57.5\%$). According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum, identified with the color yellow, is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5 , in fact, the expected value of $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}=S_{i}^{(0)}]$ for the counties allocated to this strata has a median close to zero and the first and third quantiles of $-0.29\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum, identified with the color green, is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5 . Specifically, in these counties, the NAAQS revision reduced by $0.9\mu g/m^{3}$ the median levels pm2.5 , with first and third quantiles of $-1.46\mu g/m^{3}$ and $-0.38\mu g/m^{3}$. The associative positive stratum, identified with the color red, is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.8\mu g/m^{3}$ in the median. However, this latter stratum has the highest uncertainty for $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}\prec S_{i}^{(0)}]$, with values of the first and third quantiles equal to $0.13\mu g/m^{3}$ and $1.34\mu g/m^{3}$, respectively.
@@
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata, i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The EDE and the EAE+ have a similar increase in the median $2.7{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ of the mortality rate when environmental plans are applied. The means of these two effects are $0.8{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ and $1.3{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$, respectively. For both principal causal effects, the interquartile interval includes the zero value. This means that in counties where the NAAQS revision does not affect the level pm2.5 or where this level increases, the mortality rate is also not affected. The EAE-, in green color on the right image of Figure 2, assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5 , significantly reducing it, also reduces the median age adjusted mortality rate. Specifically, the mortality rate decreases by $11.9{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median and with an interquartile interval of $[-15.2{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}},-9.1{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}]$.
@@
Figure 2: Boxplots of the three identified strata for (left) the conditional average of the difference of the post-treatment variables, and (right) the expected associative/dissociative effects. The light-blue vertical lines show the value zero, identifying the null effects.
@@
In addition, it is our interest to characterize the three different strata (Figure F3 in the Supplementary Materials visualizes the average of observed covariates within each stratum, reported as colored lines, and compares them with the average of covariates between the full 384 counties). The associative positive stratum and the dissociative stratum are composed of rural counties with a smaller population density, where the population has lower income with respect to the mean of the overall counties, and with a small employment rate. In contrast, in countries of the associative positive stratum, the percentage of Hispanics is lower than the average, in favor of a higher percentage of whites and Asians, while the dissociative stratum is characterized by a higher percentage of Hispanics and a smaller community of black people, with a higher level of education. Furthermore, the associative positive stratum is distinguished by a high smoke rate. In contrast, the associative negative stratum identifies urban counties with a high population density. This population is characterized by higher income, but also a higher percentage of poor people and lower levels of education. The meteorological variable, such as the averages of daily temperatures, the relative average of humidity, and the dew point, appears to play an important role in the distinction of strata, that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
@@ 6 Discussion
In this paper, we proposed the confounders-aware shared atoms Bayesian mixture model to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. Our proposed idea of allowing the distribution of potential post-treatment variables to share information through the shared-atoms of the Bayesian nonparametric prior enables the accurate identification of principal strata while relaxing the standard assumptions of the principal stratification causal framework. Our proposed model data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is non-zero and depends on its observed covariates.
The performance of the proposed model is evaluated through an extensive set of Monte Carlo simulations where the performance of our proposed model is compared and contrasted with the performance of the proposed models by Schwartz et al. (2011) and Lu et al. (2023). The results demonstrate that our model consistently outperforms these alternatives, achieving lower bias, better principal strata identification, and accurate estimation of principal causal effects across various scenarios.

### v3

@@ 5.1 Data and Study Design
In our application, we investigate the repercussions of the National Ambient Air Quality Standards (NAAQS) revision for air pollution in the U.S. This revision dictated a more stringent environmental policy for the concentration of pm2.5. First, we evaluate the direct and indirect effects of the 2005 NAAQS revision on the mortality rate in the period 2010-2016, considering the variation of levels pm2.5, under the principal stratification framework. Second, we take advantage of the flexible Bayesian non-parametric mixture model for the post-treatment variable (that is, pm2.5) to understand how these effects can vary across principal strata. Third, we provide a characterization of these different identified groups.
To answer our research question, we have merged two datasets: (i) one dataset containing the information about the NAAQS designations, as well as pm2.5 and the demographic and socioeconomic characteristics in the counties in the Eastern U.S., used in the Zigler et al. (2018)’s analysis (data can be found at Zigler, 2017), and (ii) one dataset containing the information on the age-adjusted mortality rate in these counties available by the Center for Disease Control and Prevention (CDC) of U.S. (Friede et al., 1993).
@@ 5.2 Results
We apply CASBAH to the 384 counties in the Eastern US described above, including all covariates, census and meteorological variables, in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 124 counties ($32\%$ of the total counties analyzed), the associative positive stratum with 46 counties ($12\%$), and the associative negative stratum with 214 counties ($56\%$).
@@
According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum, identified with the color yellow, is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5, in fact, the expected value of $\mathbb{E}\{P_{i}(1)-P_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}\}$ for the counties allocated to this strata has a median close to zero and the $90\%$ credible intervals of $0.46\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum, identified with the color green, is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5. Specifically, in these counties, the NAAQS revision reduced by $1.09\mu g/m^{3}$ the median levels pm2.5, with $90\%$ credible intervals of of $1.32\mu g/m^{3}$ and $0.82\mu g/m^{3}$. The associative positive stratum, identified with the color red, is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.50\mu g/m^{3}$ in the median and credible intervals equal to $0.19\mu g/m^{3}$ and $0.72\mu g/m^{3}$, respectively.
@@
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata, i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The associative negative effect, in green color on the right image of Figure 2, assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5, reduce it, also reduces the median age-adjusted mortality rate. Specifically, the mortality rate decreases by $8.12{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median. The dissociative effect (in yellow) shows a decrease of $8.16{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in median of the mortality rate when environmental plans are applied, while the associative positive effects decrease in median by $17.87{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$.
@@
Figure 2: Posterior medians and $90\%$ credible intervals of the three identified strata for (left) the conditional average of the difference of the post-treatment variables, and (right) the expected associative/dissociative effects. The x-axes indicate (left) the pm2.5 variation in $\mu g/m^{3}$ and (right) the mortality variation in $^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$. The light-blue vertical lines show the value zero, identifying the null effects.
@@
In addition, it is our interest to characterize the three different strata. Figure3 visualizes the average of observed covariates within each stratum, reported as colored lines, and compares them with the average of covariates between the full 384 counties in gray. The associative positive stratum and the dissociative stratum are composed of rural counties with a higher percentage of women and Black communities, where the population has lower income relative to the mean of the overall counties and with a small employment rate. In addition, the positive associative stratum (in red) is also characterized by a lower education rate, higher Hispanic community and a higher percentage of smokers.
In contrast, the counties in the associative negative stratum (in green) are mainly composed by urban areas with a high population density, high levels of education, higher income and house values, higher rate of move, and more men. The meteorological variables, such as the averages of daily temperatures, the relative average of humidity, and the dew point, also appear to play an important role in the positive associative stratum—that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
Figure 3: Representation of the characteristics of the identified strata. Each spider plot reports in the colored area the strata-specific characteristics (the mean of the analyzed covariates) and in the gray area the collective characteristics (the mean of the covariates among all the analyzed counties in the Eastern U.S.). We can consider the gray area as the benchmark to understand how the characteristics of each stratum differ from the collective characteristics of the analyzed population.
@@ 6 Discussion
In this paper, we introduced the we proposed the Confounders-Aware SHared Atoms Bayesian Mixture Model (CASBAH) to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. An innovate feature of CASBAH is that we allow the distribution of potential post-treatment variables to share information through the shared-atoms of the Bayesian nonparametric prior. This enables a more accurate identification of the principal strata while relaxing the standard assumptions of the causal framework of the principal stratification.
CASBAH data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is nonzero and depends on its observed covariates.
The performance of the proposed model is evaluated through an extensive set of Monte Carlo simulations where the performance of CASBAH is compared and contrasted with the performance of the proposed models by Schwartz et al. (2011) and Lu et al. (2023). The results demonstrate that CASBAH consistently outperforms these alternatives, achieving lower bias, better principal strata identification, and accurate estimation of principal causal effects across various scenarios.